# Interpretable and Steerable Concept Bottleneck Sparse Autoencoders

### v1

@@ 5.2 Training CB-SAE
Step 4 (Fig. 4) We introduce a concept bottleneck autoencoder [kulkarni2025interpretable] alongside the retained SAE (Fig. 1B). Our CB-SAE consists of the retained SAE $E^{\prime}_{\text{sae}},D^{\prime}_{\text{sae}}$, a linear concept encoder $E_{\text{cb}}\in\mathbb{R}^{|\mathcal{C}|\times d}$, a linear concept decoder $D_{\text{cb}}\in\mathbb{R}^{d\times|\mathcal{C}|}$, and a non-linear activation function $\sigma_{\text{cb}}:\mathbb{R}^{|\mathcal{C}|}\to\mathbb{R}^{|\mathcal{C}|}$ similar to the SAE, where $\mathcal{C}$ is a pre-defined concept set. For an input $v\in\mathbb{R}^{d}$, the CB-SAE reconstructs $\hat{v}^{\prime}\in\mathbb{R}^{d}$ as,
@@
Concept Set Selection. Based on our motivation to support user-specified concepts, the concept set $\mathcal{C}$ can be specified by the user as a list of text-based concepts, similar to prior work [oikarinen2023labelfree, srivastava2024vlg]. However, as shown in Sec. 4, the SAE can already represent some concepts well and including them in the concept set $\mathcal{C}$ would be redundant. Hence, we only use concepts absent from the retained SAE in the CB-SAE concept set. Let $\mathcal{C}_{\text{user}}$ be the user-specified concept set, $\mathcal{C}_{\text{rsae}}\subset\mathcal{C}_{\text{user}}$ be the concepts present in the retained SAE (found using CLIP-Dissect before pruning the SAE, Fig. 4 Step 2). Then our CB-SAE concept set is given by $\mathcal{C}=\mathcal{C}_{\text{user}}\setminus\mathcal{C}_{\text{rsae}}$.
@@
Objective B: Interpretability $\mathbf{\mathcal{L}_{\text{int}}}$ (Fig. 4, Step 4B). To ensure that each CB neuron in $c$ activates for its corresponding concept in the concept set $\mathcal{C}$, we use a CLIP zero-shot classifier [radford2021learning] with $\mathcal{C}$ as the classes and obtain pseudo-ground-truth concept activations similar to prior work in CBMs [oikarinen2023labelfree, kulkarni2025interpretable]. This enables concept-label-free training of the CB-SAE, i.e. does not require explicit concept labels and also supports any arbitrary user-specified concept sets.
The zero-shot classifier $\mathcal{M}:\mathbb{X}\times\mathbb{T}^{|\mathcal{C}|}\to\mathbb{R}^{|\mathcal{C}|}$ ($\mathbb{T}$ refers to text space) takes in an image $x\in\mathbb{X}$, list of concept names $\mathcal{C}$, and predicts concept logits $\hat{y}=\mathcal{M}(x,\mathcal{C})\in\mathbb{R}^{|\mathcal{C}|}$. We use a cosine-cubed similarity loss $\mathcal{L}_{\text{int}}$ following [oikarinen2023labelfree] between the concept encoder’s predictions $c=E_{\text{cb}}(v)$ and $\hat{y}$,
@@
Here, $v=f_{l}(x)$ is the vision encoder $f$ output at layer $l$ for the same image $x$ used with $\mathcal{M}$. Further, note that we use the sparsity constraint of a top-$k$ activation function $\sigma_{\text{cb}}$ only for the decoder in Eq. (7). This is because the concept encoder $E_{\text{cb}}$ should be able to interpret all concepts in an image, but the concept decoder $D_{c}$ can only use the top-$k$ concepts for reconstruction. This is further ensured by updating only $E_{\text{cb}}$ with the interpretability objective $\mathcal{L}_{\text{int}}$. We defer more details of the cosine-cubed similarity loss [oikarinen2023labelfree] to the Appendix.
@@
Downstream
Task
Steered
Model
Method Interpretability Steerability
@@
Image + Text $\to$
Text Generation
LLaVA-1.5-7B [liu2024improved]
(CLIP-ViT-L + Vicuna-7B)
SAE [pach2025sparse] 0.154 0.517 0.198 0.203
@@
LLaVA-MORE [cocchi2025llava]
(DINOv2-L + Gemma2-9B)
SAE [pach2025sparse] 0.194 0.553 0.179 0.177
@@
Image $\to$ Image
Generation
UnCLIP [Rombach_2022_CVPR]
(CLIP-ViT-L + SD-2.1)
SAE [pach2025sparse] 0.058 0.540 0.642 0.654
@@
Objective C: Steerability $\mathbf{\mathcal{L}_{\text{st}}}$ (Fig. 4, Step 4C). Prior work on CBMs for generative models [kulkarni2025interpretable] leveraged the downstream image generation task to design explicit steerability objectives. In contrast, we propose a simple task-agnostic cyclic reconstruction objective for steerability. With this, we show in Sec. 6 that the same CB-SAE can steer two different downstream tasks: image-to-image generation and image-text-to-text generation.
@@
Instead of loss weighting hyperparameters, we train by alternately minimizing the objectives via separate Adam optimizers which adaptively scale weight updates [kundu2019bihmp].
@@ 6.1 Setup
Baseline SAE and CB-SAE. We follow pach2025sparse and train a Matryoshka Batch Top-$k$ SAE [zaigrajew2025interpreting] with expansion factor $\frac{\omega}{d}=64$ as the baseline SAE on the ImageNet-1k [deng2009imagenet] dataset. Our CB-SAE is also trained on the same intermediate activations as the baseline SAE for a fair comparison. We retain $\omega-M=30$k neurons in the SAE pruning and use a top-$k$ function as $\sigma_{\text{cb}}$ with $k=5$ in our CB-SAE. We use the VLG-CBM ImageNet concept set [srivastava2024vlg, oikarinen2023labelfree] for the CB neurons. In our training, we use a CLIP-ViT-B/16 [radford2021learning] model for obtaining the pseudo-ground-truth concept activations.
Evaluation Metrics. To evaluate interpretability, we use the CLIP-Dissect interpretability score introduced in Sec. 3 and the monosemanticity score from [pach2025sparse] using the ImageNet validation set. To ensure a fair evaluation, we use a stronger CLIP-ViT-L/14 model (w.r.t. smaller ViT-B/16 used for training CB-SAE). To evaluate steerability, we use our proposed steerability score (Sec. 3). Concretely, we evaluate the steerability of each CB/SAE neuron in two ways:
  • Unit Vector: The selected neuron is activated to a high value $\alpha=50$ (as in [pach2025sparse]) & all other neurons are set to 0.
  • White Image: The selected neuron is activated to a high value $\alpha=50$ and all other neurons have the values predicted when using an empty white image (following [pach2025sparse]) as input, instead of 0 like in unit vector steering.
The interpretability and steerability scores of individual neurons are averaged to obtain the overall scores. For experiments where the steered output is text, we compare the similarity between the steered text and the CLIP-Dissect assigned concept for the selected neuron in a sentence transformer embedding space (as in Sec. 3). For experiments where steered output is an image, we compute the average similarity between the steered image and top-16 highly activating images for the selected neuron in the DINOv2 [oquab2024dinov2] embedding space. This is because the diffusion model being steered (UnCLIP [Rombach_2022_CVPR]) may rarely return partially or completely noisy images after steering, which cannot not be properly evaluated with an image-text similarity score (e.g. CLIP) that expects clean images. All metrics are normalized in 0-1 range and higher values indicate better performance.
@@ 6.2 Quantitative Comparison
In Table 1, we compare our CB-SAE with the baseline SAE [pach2025sparse] across several downstream models as well as tasks. For two different variants of the (image,text)-to-text LLaVA [liu2024improved, cocchi2025llava] model, our CB-SAE demonstrates consistent gains over the SAE baseline across both interpretability (avg. +33.0% for CLIP+Vicuna and avg. +29.0% for DINOv2+Gemma) and steerability metrics (avg. +27.5% for CLIP+Vicuna and avg. +14.0% for DINOv2+Gemma). Similarly, for the image-to-image UnCLIP [Rombach_2022_CVPR] generative model, our CB-SAE outperforms the SAE baseline (avg. +34.3% interpretability and avg. +2.1% steerability). To the best of our knowledge, we are the first to show that an SAE (and CB-SAE) trained with the same method can be used to steer different downstream tasks.
@@ 6.3 Analysis of our CB-SAE
Qualitative examples of steering. We report qualitative examples of steering LLaVA and UnCLIP with discarded SAE neurons, retained SAE neurons, and CB neurons in Fig. 7. We find discarded SAE neurons to be worse in steering than retained SAE neurons, which is expected as pruning uses the steerability score. Specifically, in the image-to-image generator UnCLIP [Rombach_2022_CVPR], we find steering the CB neurons produce higher quality images compared to the SAE neurons which often produce noisy images, likely due to the explicit concept supervision in CB-SAE. We also highlight some failure cases and partially correct steering for all neurons. We provide white image steering examples for UnCLIP in the Appendix due to space constraints.
@@ 7 Conclusion
In this work, we made the first attempt to unify two complementary paradigms- SAEs for unsupervised concept discovery and CBM for interpretable control-into a single unified framework, CB-SAE. Motivated by insights derived from our comprehensive analysis of SAEs in LVLMs, we first pruned the low-utility neurons and their corresponding weights in the SAE. We then introduced a light-weight CB module trained alongside with the frozen, retained SAE using three principled objectives. Through systematic evaluation across two different downstream generation settings- vision-language assistance (LLaVA) and image generation (UnCLIP) tasks, we demonstrate that CB-SAE consistently improves both interpretability and steerability while enabling explicit user-specified concept control.
We also acknowledge that the efficacy of our approach depends on the reliability of CLIP-Dissect in assigning accurate neuron-level concepts; however, continued advances in vision-language models are likely to enhance its performance. Extending and exploring hybrid approaches that combine the strengths of other unsupervised concept discovery methods such as transcoders [paulo2025transcoders] with user-specified concept control methods constitute our future work.
@@
  • Section A: Implementation Details
@@
      Extended analysis (Sec. B.3, Table 3, 4, Fig. 9)
@@
      Extended qualitative results (Sec. B.4, Fig. 10)
Appendix A Implementation Details
Intepretability Score. We define our CLIP-Dissect-based interpretability score as the similarity score obtained from CLIP-Dissect, averaged across all SAE/CB-SAE neurons.
CLIP-Dissect [oikarinen2023clip]. Consider a probing dataset of $N$ images $\mathcal{D}=\{x_{i}\in\mathbb{X}\}_{i=1}^{N}$ where $\mathbb{X}$ is the space of images, a concept set $\mathcal{C}=\{c_{k}\}_{k=1}^{M}$ with $M$ concepts in text form, and let layer $l$ of model $f$ being explained be denoted by $f_{l}$. CLIP-Dissect uses the probing set and a multimodal model, e.g. CLIP [radford2021learning] with an image and text encoder $E_{I},E_{T}$ to identify concepts from $\mathcal{C}$ for individual neurons at the output of $f_{l}$.
The probing set $\mathcal{D}$ is passed through the CLIP image encoder $E_{I}$ to obtain corresponding set of image embeddings $\{A_{i}=E_{I}(x_{i})\}_{i=1}^{N}$. The concept set is passed through the CLIP text encoder $E_{T}$ to obtain text embeddings $\{E_{T}(c_{k})\}_{k=1}^{M}$. Next, a matrix $P\in\mathbb{R}^{N\times M}$ is computed as the inner product of the image-text embeddings with entries $P_{ik}=A_{i}^{\top}E_{T}(c_{k})$, as CLIP image and text encoders have the same embedding dimensions. The layer $l$ activations of a neuron $j$ for the same probing set are denoted by $q_{j}=[f_{l}(x_{1})_{j},f_{l}(x_{2})_{j},\cdots,f_{l}(x_{N})_{j}]$. Finally, each neuron $j$ can be identified to have the concept $\arg\max_{k}\text{sim}(P_{:,k},q_{j})$ where $P_{:,k}$ is the $k^{\text{th}}$ column of $P$. In other words, we compare each neuron’s activations over the probing set with the corresponding activations of the CLIP model for each concept, and select the concept with the highest similarity. The maximum similarity itself (averaged across all neurons) is used as our interpretability score. The similarity function sim is soft weighted pointwise mutual information (soft-WPMI) following [oikarinen2023clip]. Please refer to the original paper [oikarinen2023clip] for more details.
Cosine-cubed similarity loss [oikarinen2023labelfree] $\mathcal{L}_{\text{int}}$. As discussed in Sec. 5.2 (main paper), we use a cosine-cubed similarity loss $\mathcal{L}_{\text{int}}$ to train the CB encoder $E_{\text{cb}}$ to produce concept predictions $c$ that match with CLIP zero-shot classifier predictions $\hat{y}$ for the same concept set $\mathcal{C}$. Concretely,

### v2

@@ 5.2 Training CB-SAE
Step 4 (Fig. 4) We introduce a concept bottleneck autoencoder [25] alongside the retained SAE (Fig. 1B). Our CB-SAE consists of the retained SAE $E^{\prime}_{\text{sae}},D^{\prime}_{\text{sae}}$, a linear concept encoder $E_{\text{cb}}\in\mathbb{R}^{|\mathcal{C}|\times d}$, a linear concept decoder $D_{\text{cb}}\in\mathbb{R}^{d\times|\mathcal{C}|}$, and a non-linear activation function $\sigma_{\text{cb}}:\mathbb{R}^{|\mathcal{C}|}\to\mathbb{R}^{|\mathcal{C}|}$ similar to the SAE, where $\mathcal{C}$ is a pre-defined concept set. For an input $v\in\mathbb{R}^{d}$, the CB-SAE reconstructs $\hat{v}^{\prime}\in\mathbb{R}^{d}$ as,
@@
Concept Set Selection. Based on our motivation to support user-specified concepts, the concept set $\mathcal{C}$ can be specified by the user as a list of text-based concepts, similar to prior work [34, 47]. However, as shown in Sec. 4, the SAE can already represent some concepts well and including them in the concept set $\mathcal{C}$ would be redundant. Hence, we only use concepts absent from the retained SAE in the CB-SAE concept set. Let $\mathcal{C}_{\text{user}}$ be the user-specified concept set, $\mathcal{C}_{\text{rsae}}\subset\mathcal{C}_{\text{user}}$ be the concepts present in the retained SAE (found using CLIP-Dissect before pruning the SAE, Fig. 4 Step 2). Then our CB-SAE concept set is given by $\mathcal{C}=\mathcal{C}_{\text{user}}\setminus\mathcal{C}_{\text{rsae}}$.
@@
Objective B: Interpretability $\mathbf{\mathcal{L}_{\text{int}}}$ (Fig. 4, Step 4B). To ensure that each CB neuron in $c$ activates for its corresponding concept in the concept set $\mathcal{C}$, we use a CLIP zero-shot classifier [42] with $\mathcal{C}$ as the classes and obtain pseudo-ground-truth concept activations similar to prior work in CBMs [34, 25]. This enables concept-label-free training of the CB-SAE, i.e. does not require explicit concept labels and also supports any arbitrary user-specified concept sets.
The zero-shot classifier $\mathcal{M}:\mathbb{X}\times\mathbb{T}^{|\mathcal{C}|}\to\mathbb{R}^{|\mathcal{C}|}$ ($\mathbb{T}$ refers to text space) takes in an image $x\in\mathbb{X}$, list of concept names $\mathcal{C}$, and predicts concept logits $\hat{y}=\mathcal{M}(x,\mathcal{C})\in\mathbb{R}^{|\mathcal{C}|}$. We use a cosine-cubed similarity loss $\mathcal{L}_{\text{int}}$ following [34] between the concept encoder’s predictions $c=E_{\text{cb}}(v)$ and $\hat{y}$,
@@
Here, $v=f_{l}(x)$ is the vision encoder $f$ output at layer $l$ for the same image $x$ used with $\mathcal{M}$. Further, note that we use the sparsity constraint of a top-$k$ activation function $\sigma_{\text{cb}}$ only for the decoder in Eq. (7). This is because the concept encoder $E_{\text{cb}}$ should be able to interpret all concepts in an image, but the concept decoder $D_{c}$ can only use the top-$k$ concepts for reconstruction. This is further ensured by updating only $E_{\text{cb}}$ with the interpretability objective $\mathcal{L}_{\text{int}}$. We defer more details of the cosine-cubed similarity loss [34] to the Appendix.
@@
Downstream Task Steered Model Method Interpretability Steerability
@@
Image + Text $\to$ Text Generation LLaVA-1.5-7B [28] (CLIP-ViT-L + Vicuna-7B) SAE [37] 0.154 0.517 0.198 0.203
@@
LLaVA-MORE [12] (DINOv2-L + Gemma2-9B) SAE [37] 0.194 0.553 0.179 0.177
@@
Image $\to$ Image Generation UnCLIP [46] (CLIP-ViT-L + SD-2.1) SAE [37] 0.058 0.540 0.642 0.654
@@
Objective C: Steerability $\mathbf{\mathcal{L}_{\text{st}}}$ (Fig. 4, Step 4C). Prior work on CBMs for generative models [25] leveraged the downstream image generation task to design explicit steerability objectives. In contrast, we propose a simple task-agnostic cyclic reconstruction objective for steerability. With this, we show in Sec. 6 that the same CB-SAE can steer two different downstream tasks: image-to-image generation and image-text-to-text generation.
@@
Instead of loss weighting hyperparameters, we train by alternately minimizing the objectives via separate Adam optimizers which adaptively scale weight updates [26].
@@ 6.1 Setup
Baseline SAE and CB-SAE. We follow Pach et al. [37] and train a Matryoshka Batch Top-$k$ SAE [63] with expansion factor $\frac{\omega}{d}=64$ as the baseline SAE on the ImageNet-1k [14] dataset. Our CB-SAE is also trained on the same intermediate activations as the baseline SAE for a fair comparison. We retain $\omega-M=30$k neurons in the SAE pruning and use a top-$k$ function as $\sigma_{\text{cb}}$ with $k=5$ in our CB-SAE. We use the VLG-CBM ImageNet concept set [47, 34] for the CB neurons. In our training, we use a CLIP-ViT-B/16 [42] model for obtaining the pseudo-ground-truth concept activations.
Evaluation Metrics. To evaluate interpretability, we use the CLIP-Dissect interpretability score introduced in Sec. 3 and the monosemanticity score from [37] using the ImageNet validation set. To ensure a fair evaluation, we use a stronger CLIP-ViT-L/14 model (w.r.t. smaller ViT-B/16 used for training CB-SAE). To evaluate steerability, we use our proposed steerability score (Sec. 3). Concretely, we evaluate the steerability of each CB/SAE neuron in two ways:
  • Unit Vector: The selected neuron is activated to a high value $\alpha=50$ (as in [37]) & all other neurons are set to 0.
  • White Image: The selected neuron is activated to a high value $\alpha=50$ and all other neurons have the values predicted when using an empty white image (following [37]) as input, instead of 0 like in unit vector steering.
The interpretability and steerability scores of individual neurons are averaged to obtain the overall scores. For experiments where the steered output is text, we compare the similarity between the steered text and the CLIP-Dissect assigned concept for the selected neuron in a sentence transformer embedding space (as in Sec. 3). For experiments where steered output is an image, we compute the average similarity between the steered image and top-16 highly activating images for the selected neuron in the DINOv2 [36] embedding space. This is because the diffusion model being steered (UnCLIP [46]) may rarely return partially or completely noisy images after steering, which cannot be properly evaluated with an image-text similarity score (e.g. CLIP) that expects clean images. All metrics are normalized in 0-1 range and higher values indicate better performance.
@@ 6.2 Quantitative Comparison
In Table 1, we compare our CB-SAE with the baseline SAE [37] across several downstream models as well as tasks. For two different variants of the (image,text)-to-text LLaVA [28, 12] model, our CB-SAE demonstrates consistent gains over the SAE baseline across both interpretability (avg. +33.0% for CLIP+Vicuna and avg. +29.0% for DINOv2+Gemma) and steerability metrics (avg. +27.5% for CLIP+Vicuna and avg. +14.0% for DINOv2+Gemma). Similarly, for the image-to-image UnCLIP [46] generative model, our CB-SAE outperforms the SAE baseline (avg. +34.3% interpretability and avg. +2.1% steerability). To the best of our knowledge, we are the first to show that an SAE (and CB-SAE) trained with the same method can be used to steer different downstream tasks.
@@ 6.3 Analysis of our CB-SAE
Qualitative examples of steering. We report qualitative examples of steering LLaVA and UnCLIP with discarded SAE neurons, retained SAE neurons, and CB neurons in Fig. 7. We find discarded SAE neurons to be worse in steering than retained SAE neurons, which is expected as pruning uses the steerability score. Specifically, in the image-to-image generator UnCLIP [46], we find steering the CB neurons produce higher quality images compared to the SAE neurons which often produce noisy images, likely due to the explicit concept supervision in CB-SAE. We also highlight some failure cases and partially correct steering for all neurons.
@@ 7 Conclusion
In this work, we made the first attempt to unify two complementary paradigms - SAEs for unsupervised concept discovery and CBM for interpretable concept control - into a single unified framework, CB-SAE. Motivated by insights derived from our comprehensive analysis of SAEs in LVLMs, we first pruned the low-utility neurons in the SAE. We then introduced a light-weight CB module trained alongside the frozen, retained SAE using three principled objectives. Across vision-language (LLaVA) and image generation (UnCLIP) tasks, CB-SAE consistently improves interpretability and steerability, while enabling user-specified concept control.
Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by the Lawrence Livermore National Laboratory under Contract No. DE-AC52-07NA27344, Lawrence Livermore National Security, LLC. and was supported by the LLNL-LDRD Program under Project No. 25-SI-001 and DOE ECRP 51917/SCW1885. LLNL-CONF-2013863. A. Kulkarni and T-W. Weng are also partially supported by National Science Foundation under Grant No. 2313105, 2430539, Hellman Fellowship, ARL Award, and Intel Rising Star Faculty Award.
@@ Acknowledgements
  • Section A: Limitations and Future Work
  • Section B: Implementation Details
@@
      Extended analysis (Sec. C.3, Table 3, 4, 5, 6, Fig. 9)
@@
      Extended qualitative results (Sec. C.4, Fig. 10)
Appendix A Limitations and Future Work
We acknowledge that the efficacy of our approach depends on the reliability of CLIP-Dissect in assigning accurate neuron-level concepts; however, continued advances in vision-language models are likely to enhance its performance. Extending and exploring hybrid approaches that combine the strengths of other unsupervised concept discovery methods such as transcoders [41] with user-specified concept control methods constitute our future work.
Recent work on LLM SAEs [10] showed “feature-splitting” where hierarchical features split into fine-grained features (e.g. “math” may be represented in the SAE by “algebra” and “geometry” neurons). It will be interesting to investigate if this phenomenon is connected to SAEs having limited concept coverage that we highlighted in this paper.
Appendix B Implementation Details
CLIP-Dissect [35]. Consider a probing dataset of $N$ images $\mathcal{D}=\{x_{i}\in\mathbb{X}\}_{i=1}^{N}$ where $\mathbb{X}$ is the space of images, a concept set $\mathcal{C}=\{c_{k}\}_{k=1}^{M}$ with $M$ concepts in text form, and let layer $l$ of model $f$ being explained be denoted by $f_{l}$. CLIP-Dissect uses the probing set and a multimodal model, e.g. CLIP [42] with an image and text encoder $E_{I},E_{T}$ to identify concepts from $\mathcal{C}$ for individual neurons at the output of $f_{l}$.
The probing set $\mathcal{D}$ is passed through the CLIP image encoder $E_{I}$ to obtain corresponding set of image embeddings $\{A_{i}=E_{I}(x_{i})\}_{i=1}^{N}$. The concept set is passed through the CLIP text encoder $E_{T}$ to obtain text embeddings $\{E_{T}(c_{k})\}_{k=1}^{M}$. Next, a matrix $P\in\mathbb{R}^{N\times M}$ is computed as the inner product of the image-text embeddings with entries $P_{ik}=A_{i}^{\top}E_{T}(c_{k})$, as CLIP image and text encoders have the same embedding dimensions. The layer $l$ activations of a neuron $j$ for the same probing set are denoted by $q_{j}=[f_{l}(x_{1})_{j},f_{l}(x_{2})_{j},\cdots,f_{l}(x_{N})_{j}]$. Finally, each neuron $j$ can be identified to have the concept $\arg\max_{k}\text{sim}(P_{:,k},q_{j})$ where $P_{:,k}$ is the $k^{\text{th}}$ column of $P$. In other words, we compare each neuron’s activations over the probing set with the corresponding activations of the CLIP model for each concept, and select the concept with the highest similarity. The maximum similarity itself (averaged across all neurons) is used as our interpretability score. The similarity function sim is soft weighted pointwise mutual information (soft-WPMI) following [35]. Please refer to the original paper [35] for more details.
Cosine-cubed similarity loss [34] $\mathcal{L}_{\text{int}}$. As discussed in Sec. 5.2 (main paper), we use a cosine-cubed similarity loss $\mathcal{L}_{\text{int}}$ to train the CB encoder $E_{\text{cb}}$ to produce concept predictions $c$ that match with CLIP zero-shot classifier predictions $\hat{y}$ for the same concept set $\mathcal{C}$. Concretely,